# From AI for Science to Agentic Science: A Survey on Autonomous Scientific Discovery

### v1

@@ 1 Introduction
Scientific discovery is experiencing a transformative shift, driven by the rapid evolution of artificial intelligence (AI) from specialized tools to collaborative research partners. This progression marks a pivotal stage in the AI for Science paradigm, where AI systems have moved from acting as computational oracles for targeted tasks [121, 306, 339, 42, 87, 281, 351] toward the emergence of Agentic Science (Figure 1[229, 220, 81, 274, 157]. Agentic Science denotes a specific stage within AI for Science evolution-corresponding primarily to Level 3 (Full Agentic Discovery) with precursors at Level 2 (Partial Agentic Discovery) in Figure 1. In this stage, AI operates as an autonomous scientific agent capable of formulating hypotheses, designing and executing experiments, interpreting results, and iteratively refining theories with reduced dependence on human guidance [229, 22]. This advancement is enabled by integrated platforms such as Intern-Discovery, which grants access to diverse AI agents and datasets, and by multimodal models like Intern-S1111https://github.com/InternLM/Intern-S1, which demonstrate deep scientific reasoning.
This transformation is fueled by recent breakthroughs in foundational models, particularly large language models (LLMs) [82, 256, 351], which provide unprecedented capabilities in natural language understanding, complex reasoning, and tool use [245, 348, 323, 338, 337]. These capabilities have enabled the development of AI agents that transcend static learning pipelines, acting instead as dynamic, goal-driven entities navigating the scientific method [311, 89, 178, 358]. From hypothesis generation [307, 209] to autonomous experimentation [22, 317] and synthetic dataset creation [150], these agents demonstrate emergent behaviors once considered exclusively human.
@@ 2.1.1 Level 1: AI as a Computational Oracle (Expert Tools)
This paradigm has led to significant advances across the natural sciences. In life sciences, AI has revolutionized genomics [45, 10, 191], proteomics [2, 121, 219, 283, 339, 214, 90, 340, 122], and single-cell analysis [87, 42]. In chemistry, models like MolGPT [13], ChemLLM [328], and ChemMLLM [252] have accelerated molecular design and property prediction [109]. Materials science has benefited from AI-driven discovery platforms [16, 302, 300, 301, 324]. In physics and astronomy, AI tools are used for tasks ranging from modeling quantum systems [201, 342, 270] and detecting phase transitions [80, 29] to analyzing astronomical data [192, 62, 281] and modeling fluid dynamics [51, 354, 156].
@@ 2.3 Agentic Science: The Focus of This Survey
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=4pt, line width=1.1pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=9.5em,font=,, where level=2text width=11.5em,font=,, where level=3text width=12em,font=,, where level=4text width=50em,font=,, [Taxonomy, ver [Core Abilities and
Challenges of
Scientific Agents [Planning and
Reasoning Engines [e.g., CoT [130], self-consistency [275], debate [255], ToT [97], MCTS [83], ReAct [310], etc., leaf3, text width=44em] ] [Tool Use and Integration [e.g., Code interpreters [78], search engines [109], ChemCrow [24], CRISPR-GPT [103],
MuJoCo [262], docking tools [60], etc., leaf3, text width=44em] ] [Memory Mechanism [e.g., Reflexion [233], Voyager [268], RAG [141], DrugAgent [108], GraphRAG [54], etc., leaf3, text width=44em] ] [Collaboration
between Agents [e.g., MetaGPT [93], Coscientist [22], AutoGen [286], ReConcile [36], DyLAN [166], etc., leaf3, text width=44em] ] [Optimization
and Evolution [e.g., SELF-REFINE [179], CRITIC [79], RL with self-reward [319], KnowAgent [357], CAMEL [143],
debate [52], etc., leaf3, text width=44em] ] ] [Core Process
Challenges of
Agentic Science [Fully Autonomous
Research Pipeline [e.g., NovelSeek [257], The AI Scientist [169], X-Master [32], Biomni [104], Coscientist [22],
LLM-RDF [223], MatPilot [194], PROTEUS [50], The Virtual Lab [248], AI co-scientist [76],
VirSci [242], MAPS [334], DORA [189], MDAgents [128], DiscoveryWorld [112], STELLA [120],
AgentRxiv [227], Agent Hospital [145], ResearchAgent [11], Agent Laboratory [228],
OriGene [344], Dolphin [317], AlphaEvolve [196], Conversational Health Agents [1], etc., leaf, text width=44em] ] ] [Agentic Life
Science Research [General Frameworks
and Methodologies [e.g., Biomni [104], STELLA [120], BioResearcher [171], PiFlow [208], Empowering BD [64], etc., leaf4, text width=44em] ] [Genomics, Transcriptomics,
and Multi-Omics Analysis [e.g., BIA [294], CellAgent [291], TAIS [158], CRISPR-GPT [103], SpatialAgent [269],
PhenoGraph [195],BioAgents [186], BioMaster [243], TransAgent [329], CompBioAgent [330],
PerTurboAgent [88], PROTEUS [50, 215], CellVoyager [5], AstroAgents [225],
BioDiscoveryAgent [222], OminiCellAgent [102], etc., leaf4, text width=44em] ] [Protein Science and
Engineering [e.g., ProtAgents [68], Sparks [73], etc., leaf4, text width=44em] ] [Drug and Therapeutic
Discovery [e.g., The Virtual Lab [248], OriGene [344], LLM Agent for DD [197], TxAgent [65], Robin [74],
DrugAgent [162], LIDDIA [9], PharmAgents [59], CLADD [140], Tippy [55], ACEGEN [23],
AI Co-scientist [76], Exploring Modularity [265], DO Challenge [235], PrimeGen [276],etc., leaf4, text width=44em] ] ] [Agentic Chemistry
Research [General Frameworks
and Methodologies [e.g., ChemCrow [25], ChemAgents [237], MOOSE-Chem [309], MOOSE-Chem3 [163],
ChemMiner [38], Eunomia [6], ChemAgent [285], ChemHAS [152], ChemToolAgent [315]
, Chemagent [253], LabUtopia [147], CACTUS [185], GVIM [176], MT-Mol [126], CSstep [34],
CRAG-MoW [28], etc., leaf2, text width=44em] ] [Organic Synthesis and
Reaction Optimization [e.g., Coscientist [22], LLM-RDF [224], Chemist-X [37], ORGANA [47], Dai et al. [44],
Strieth-Kalthoff et al. [241], AutoChemSchematic AI [240], etc., leaf2, text width=44em] ] [Generative Chemistry and
Molecular Design [e.g., ChatMOF [123], MOFGen [107], OSDA Agent [100], ChemReasoner [239],
Horwood & Noutahi [94], etc., leaf2, text width=44em] ] [Computational and
Quantum Chemistry [e.g., El Agente Q [359], Aitomia [96], ChemGraph [205], xChemAgents [206], etc., leaf2, text width=44em] ] ] [Agentic Materials
Research [General Frameworks and
Automated Workflows [e.g., AILA [181], Foam-Agent [320], ChemGraph [205], MechAgents [193], MatPilot [194],
LLMatDesign [115], MAPPS [349], LLaMP [40], HoneyComb [333], Bazgir et al. [17], PiFlow [208],
dZiner [7], Kumbhar et al. [134], etc., leaf5, text width=44em] ] [Structural and
Functional Materials [e.g., AtomAgents [67, 71], Ghafarollahi et al. [69], metaAgent [98], CrossMatAgent [260],
Lu et al. [170], etc., leaf5, text width=44em] ] [Advanced and
Quantum Materials [e.g., SciAgents [72], PriM [136], TopoMAS [325], etc., leaf5, text width=44em] ] ] [Agentic Physics and
Astronomy
Research [General Frameworks and
Methodologies [e.g., MoRA [111], LP-COMDA [160], LLMSat [182], CosmoAgent [298], etc., leaf6, text width=44em] ] [Astronomy and Cosmology [e.g., StarWhisper [267], mephisto [246], AI Agents [132], AI Cosmologist [188],
MAS-Cosmology [139], SimAgents [341], etc., leaf6, text width=44em] ] [Computational Mechanics
and Fluid Dynamics [e.g., OpenFOAMGPT [202], OpenFOAMGPT 2.0 [56], LLM-Agent [159], MechAgents [193],
AutoGen-FEM [259], etc., leaf6, text width=44em] ] [Quantum Computing [e.g., k-agents [30], etc., leaf6, text width=44em] ] ] ]
@@ 5.1 General Frameworks and Methodologies
STELLA [120] is a self-evolving AI agent designed to overcome the limitations of static toolsets. Its method is a multi-agent architecture featuring two core adaptive mechanisms: an evolving Template Library for reasoning strategies and a dynamic Tool Ocean that expands as a dedicated agent autonomously discovers and integrates new bioinformatics tools. This design allows STELLA to learn from experience; its results show that its accuracy on challenging biomedical benchmarks systematically improves with increased trials, outperforming leading models. Biomni [104] is presented as a general-purpose biomedical AI agent designed for flexibility across a wide array of tasks. Its method is based on decomposing complex user queries into multi-step plans and executing them by dynamically selecting from an expanding set of tools. BioResearcher [171] is another end-to-end automated system that employs a modular, multi-agent architecture for search, literature processing, experimental design, and programming. A key feature of its method is an LLM-based reviewer for in-process quality control, which enabled the system to achieve an average execution success rate of 63.07% across eight previously unmet research objectives. A more theoretical framework, PiFlow [208], recasts automated scientific discovery as a structured uncertainty reduction problem. Its information-theoretical method guides a multi-agent system’s exploration using scientific principles, which resulted in a 73.55% increase in discovery efficiency and a 94.06% enhancement in solution quality in domains including biomolecule discovery. Finally, a perspective piece envisions future "AI scientists" as collaborative agents that integrate AI models, biomedical tools, and experimental platforms [64]. The authors argue that such systems, which feature structured memory for continual learning, will empower human researchers by handling large-scale data analysis and repetitive tasks, leaving creative and strategic oversight to humans.
@@ 7.3 Automated Simulation and Characterization
Beyond simulation, agents are also being applied to automate physical experiments. AILA (Artificially Intelligent Lab Assistant) [181] is a framework that uses LLM-driven agents to automate atomic force microscopy (AFM). The work introduces AFMBench, a suite for evaluating agent performance across the scientific workflow, from experimental design to data analysis. The study found that multi-agent architectures outperform single agents and highlighted the need for rigorous benchmarking, as domain-specific knowledge (QA proficiency) did not directly translate to effective experimental control.

### v2

@@ 1 Introduction
Scientific discovery is experiencing a transformative shift, driven by the rapid evolution of artificial intelligence (AI) from specialized tools to collaborative research partners. This progression marks a pivotal stage in the AI for Science paradigm, where AI systems have moved from acting as computational oracles for targeted tasks [157, 392, 429, 61, 113, 353, 447] toward the emergence of Agentic Science (Figure 1) [288, 278, 106, 345, 203]. Agentic Science denotes a specific stage within AI for Science evolution-corresponding primarily to Level 3 (Full Agentic Discovery) with precursors at Level 2 (Partial Agentic Discovery) in Figure 1. In this stage, AI operates as an autonomous scientific agent capable of formulating hypotheses, designing and executing experiments, interpreting results, and iteratively refining theories with reduced dependence on human guidance [288, 30]. This advancement is enabled by integrated platforms such as Intern-Discovery, which grants access to diverse AI agents and datasets, and by multimodal models like Intern-S1111https://github.com/InternLM/Intern-S1, which demonstrate deep scientific reasoning.
This transformation is fueled by recent breakthroughs in foundational models, particularly large language models (LLMs) [107, 322, 447, 37, 17, 103], which provide unprecedented capabilities in natural language understanding, complex reasoning, and tool use [309, 443, 411, 428, 427]. These capabilities have enabled the development of AI agents that transcend static learning pipelines, acting instead as dynamic, goal-driven entities navigating the scientific method [397, 115, 229, 454]. From hypothesis generation [393, 267] to autonomous experimentation [30, 404] and synthetic dataset creation [194], these agents demonstrate emergent behaviors once considered exclusively human.
@@ 2.1.1 Level 1: AI as a Computational Oracle (Expert Tools)
This paradigm has led to significant advances across the natural sciences. In life sciences, AI has revolutionized genomics [65, 12, 244], proteomics [3, 157, 277, 355, 429, 272, 118, 430, 158, 432, 431], and single-cell analysis [113, 61], medicine and healthcare [128, 191, 307, 384, 129]. In chemistry, models like MolGPT [15], ChemLLM [417], and ChemMLLM [317] have accelerated molecular design and property prediction [143]. Materials science has benefited from AI-driven discovery platforms [24, 386, 382, 383, 412]. In physics and astronomy, AI tools are used for tasks ranging from modeling quantum systems [256, 434, 340] and detecting phase transitions [104, 42] to analyzing astronomical data [246, 85, 353] and modeling fluid dynamics [72, 450, 202].
@@ 2.3 Agentic Science: The Focus of This Survey
2.4 Modern Scientific Large Language Models for Agentic Science
The conceptual shift from AI as a tool to an autonomous partner is not merely theoretical; it is propelled by concrete advancements in a class of models known as Scientific Large Language Models (Sci-LLMs). These models serve as the technological bedrock for the levels of autonomy previously described. While many specialized, domain-specific models exemplify the powerful function approximators of Level 1 Computational Oracles, the recent surge in general-purpose models with advanced reasoning and tool-use capabilities provides the foundation for Level 2 and 3 Agentic Systems. This section surveys the current landscape of Sci-LLMs, illustrating the practical implementations that enable this paradigm shift.
Current scientific LLMs are mainly developed from existing general-purpose models through several common strategies. A primary approach involves fine-tuning foundational models like LLaMA on curated, science-focused instruction datasets covering disciplines such as physics, chemistry, and materials science to enhance performance on specific tasks [372, 415]. To build more robust scientific capabilities, another strategy involves large-scale pre-training or domain-adaptive pre-training on extensive scientific corpora, which include scholarly papers, textbooks, and specialized data formats like chemical formulae and protein sequences. This method strengthens a model’s core understanding of scientific principles [321, 310, 265, 18]. A third, more recent trend focuses on improving complex reasoning through test-time scaling. This is achieved by integrating techniques like Chain-of-Thought, large-scale Reinforcement Learning (RL), and massive Mixture-of-Experts (MoE) architectures, enabling models to process multimodal data and tackle multi-step scientific problems [107, 385, 322, 60, 362]. These distinct development strategies correspond to different levels of scientific autonomy: fine-tuning typically creates powerful Level 1 expert tools, while large-scale pre-training and enhanced reasoning are the key enablers for the agentic capabilities of Levels 2 and 3.
In many cases, domain-specific Sci-LLMs offer superior performance on specialized tasks. These models are constructed with well-curated datasets and training schemes tailored to a target subject. Below, we introduce recent domain-specific Sci-LLMs across eight scientific disciplines.
Life Sciences
This broad field has seen Sci-LLMs applied to multi-omics, molecular biology, and healthcare. In multi-omics, development follows two paths: foundational models trained from scratch on biological sequences (DNA, RNA, protein) to learn fundamental representations [244, 36, 280, 201, 117], and LLM-augmented systems that integrate these representations to enable conversational analysis, cross-modal translation, and controllable generation of biological sequences [199, 223, 208, 403, 63, 368, 139, 336, 375, 376, 2, 446]. Recent models aim to unify reasoning across different omics domains [363, 69]. In molecular and cellular biology, models are tuned for tasks like molecular property prediction from SMILES strings and analysis of single-cell genomic data, as well as de novo molecular design for drug discovery [212, 61, 76, 116]. In healthcare, models are often adapted from general-purpose LLMs. A common strategy is supervised fine-tuning (SFT) on medical dialogue and QA datasets [173, 31, 330, 112, 295, 346, 421]. More advanced models undergo continued pre-training on large medical corpora before SFT and reinforcement learning to improve performance and safety [357, 52, 328, 390, 370, 335]. A critical frontier is multimodal AI, where models integrate medical imaging (e.g., X-rays) with text for report generation, visual question answering (VQA), and complex diagnostic reasoning [183, 180, 214, 240, 51, 192, 289, 50, 377, 307].
Chemistry
LLMs in chemistry are designed to understand and generate information across various data modalities, including molecular structures (SMILES), reaction data, and 3D conformations. By training on specialized datasets, these models can perform a range of core chemistry tasks, such as molecular property prediction, retrosynthesis analysis, and structure-based drug design, effectively bridging text, 2D representations, and 3D geometries [417, 43, 440, 317, 152].
Materials Science
In materials science, transformer-based models are widely used for diverse applications. One approach employs encoders pre-trained on material representations like SMILES strings or scientific abstracts to predict material properties [344, 170, 168, 32]. Another utilizes decoder-only, GPT-style architectures to generate novel molecules and crystal structures with desired features [213, 21, 16, 80, 9]. More specialized models are fine-tuned for specific applications such as predicting reaction equations, assessing crystal synthesizability, and modeling material mechanics, often integrating retrieval-augmented generation (RAG) with domain knowledge graphs [19, 253, 301, 40].
Physics and Astronomy
In physics, Sci-LLMs are evolving into interactive tools for scientific workflows. These models integrate language processing with physics engines and visual modules to perform tasks such as estimating physical parameters from visual data, learning solution operators for partial differential equations (PDEs), and serving as expert knowledge retrieval systems for high-energy physics [57, 122, 436, 261]. Astronomy-specific models are typically built upon general architectures like LLaMA and adapted through continual pre-training on astronomy literature (e.g., arXiv abstracts) and fine-tuning on domain-specific tasks [142, 257, 439, 186, 239, 297, 105]. These models enhance text understanding and generation for astronomical topics and, in their multimodal variants, integrate astronomical images to perform tasks like describing galaxy images from visual data [245, 409, 156].
Taxonomy
Core Abilities and Challenges of Scientific Agents Planning and Reasoning Engines e.g., CoT [166], self-consistency [347], debate [320], ToT [127], MCTS [108], ReAct [396], etc. Tool Use and Integration e.g., Code interpreters [101], search engines [143], ChemCrow [34], CRISPR-GPT [135], MuJoCo [329], docking tools [83], etc. Memory Mechanism e.g., Reflexion [293], Voyager [338], RAG [181], DrugAgent [141], GraphRAG [75], etc. Collaboration between Agents e.g., MetaGPT [123], Coscientist [30], AutoGen [359], ReConcile [53], DyLAN [216], etc. Optimization and Evolution e.g., SELF-REFINE [230], CRITIC [102], RL with self-reward [406], KnowAgent [453], CAMEL [184], debate [73], etc. Core Process Challenges of Agentic Science Fully Autonomous Research Pipeline e.g., NovelSeek [323], The AI Scientist [219], X-Master [46], Biomni [136], Coscientist [30], LLM-RDF [282], MatPilot [248], PROTEUS [71], The Virtual Lab [313], AI co-scientist [99], VirSci [305], MAPS [424], DORA [242], MDAgents [164], DiscoveryWorld [146], STELLA [155], AgentRxiv [286], Agent Hospital [187], ResearchAgent [13], Agent Laboratory [287], OriGene [437], Dolphin [404], AlphaEvolve [250], Conversational Health Agents [1], etc. Agentic Life Science Research General Frameworks and Methodologies e.g., Biomni [136], STELLA [155], BioResearcher [221], PiFlow [266], Empowering BD [87], etc. Genomics, Transcriptomics, and Multi-Omics Analysis e.g., BIA [373], CellAgent [367], TAIS [204], CRISPR-GPT [135], SpatialAgent [339], PhenoGraph [249],BioAgents [237], BioMaster [306], TransAgent [418], CompBioAgent [419], PerTurboAgent [114], PROTEUS [71, 273], CellVoyager [6], AstroAgents [284], BioDiscoveryAgent [281], OminiCellAgent [134], etc. Protein Science and Engineering e.g., ProtAgents [91], Sparks [96], etc. Drug and Therapeutic Discovery e.g., The Virtual Lab [313], OriGene [437], LLM Agent for DD [251], TxAgent [88], Robin [97], DrugAgent [209], LIDDIA [11], PharmAgents [82], CLADD [179], Tippy [77], ACEGEN [33], AI Co-scientist [99], Exploring Modularity [333], DO Challenge [296], PrimeGen [348],etc. Agentic Chemistry Research General Frameworks and Methodologies e.g., ChemCrow [35], ChemAgents [299], MOOSE-Chem [395], MOOSE-Chem3 [210], ChemMiner [55], Eunomia [7], ChemAgent [358], ChemHAS [196], ChemToolAgent [401] , Chemagent [318], LabUtopia [189], CACTUS [236], GVIM [227], MT-Mol [162], CSstep [48], CRAG-MoW [41], etc. Organic Synthesis and Reaction Optimization e.g., Coscientist [30], LLM-RDF [283], Chemist-X [54], ORGANA [67], Dai et al. [64], Strieth-Kalthoff et al. [304], AutoChemSchematic AI [303], etc. Generative Chemistry and Molecular Design e.g., ChatMOF [159], MOFGen [140], OSDA Agent [132], ChemReasoner [302], Horwood & Noutahi [124], etc. Computational and Quantum Chemistry e.g., El Agente Q [455], Aitomia [126], ChemGraph [262], xChemAgents [263], etc. Agentic Materials Research General Frameworks and Automated Workflows e.g., AILA [232], Foam-Agent [407], ChemGraph [262], MechAgents [247], MatPilot [248], LLMatDesign [149], MAPPS [444], LLaMP [58], HoneyComb [423], Bazgir et al. [25], PiFlow [266], dZiner [8], Kumbhar et al. [172], etc. Structural and Functional Materials e.g., AtomAgents [90, 94], Ghafarollahi et al. [92], metaAgent [130], CrossMatAgent [326], Lu et al. [220], etc. Advanced and Quantum Materials e.g., SciAgents [95], PriM [175], TopoMAS [413], etc. Agentic Physics and Astronomy Research General Frameworks and Methodologies e.g., MoRA [145], LP-COMDA [206], LLMSat [233], CosmoAgent [380], etc. Astronomy and Cosmology e.g., StarWhisper [337], mephisto [311], AI Agents [169], AI Cosmologist [241], MAS-Cosmology [178], SimAgents [433], etc. Computational Mechanics and Fluid Dynamics e.g., OpenFOAMGPT [258], OpenFOAMGPT 2.0 [78], LLM-Agent [205], MechAgents [247], AutoGen-FEM [325], etc. Quantum Computing e.g., k-agents [44], etc.
@@ 5.1 General Frameworks and Methodologies
STELLA [155] is a self-evolving AI agent designed to overcome the limitations of static toolsets. Its method is a multi-agent architecture featuring two core adaptive mechanisms: an evolving Template Library for reasoning strategies and a dynamic Tool Ocean that expands as a dedicated agent autonomously discovers and integrates new bioinformatics tools. This design allows STELLA to learn from experience; its results show that its accuracy on challenging biomedical benchmarks systematically improves with increased trials, outperforming leading models. Biomni [136] is presented as a general-purpose biomedical AI agent designed for flexibility across a wide array of tasks. Its method is based on decomposing complex user queries into multi-step plans and executing them by dynamically selecting from an expanding set of tools. m-KAILIN [366] is presented as a knowledge-driven agentic framework for biomedical corpus distillation, designed to enhance large language model training. Its method is based on a multi-agent collaboration architecture guided by the MeSH knowledge hierarchy, where specialized agents autonomously generate, evaluate, and refine question–answer pairs from scientific literature to produce high-quality, ontology-aligned datasets for biomedical LLMs. BioResearcher [221] is another end-to-end automated system that employs a modular, multi-agent architecture for search, literature processing, experimental design, and programming. A key feature of its method is an LLM-based reviewer for in-process quality control, which enabled the system to achieve an average execution success rate of 63.07% across eight previously unmet research objectives. A more theoretical framework, PiFlow [266], recasts automated scientific discovery as a structured uncertainty reduction problem. Its information-theoretical method guides a multi-agent system’s exploration using scientific principles, which resulted in a 73.55% increase in discovery efficiency and a 94.06% enhancement in solution quality in domains including biomolecule discovery. Finally, a perspective piece envisions future "AI scientists" as collaborative agents that integrate AI models, biomedical tools, and experimental platforms [87]. The authors argue that such systems, which feature structured memory for continual learning, will empower human researchers by handling large-scale data analysis and repetitive tasks, leaving creative and strategic oversight to humans.
@@ 7.3 Automated Simulation and Characterization
Beyond simulation, agents are also being applied to automate physical experiments. AILA (Artificially Intelligent Lab Assistant) [232] is a framework that uses LLM-driven agents to automate atomic force microscopy (AFM). The work introduces AFMBench, a suite for evaluating agent performance across the scientific workflow, from experimental design to data analysis. The study found that multi-agent architectures outperform single agents and highlighted the need for rigorous benchmarking, as domain-specific knowledge (QA proficiency) did not directly translate to effective experimental control.The AutoMat [391] framework was developed as an agentic AI system that autonomously reconstructs atomic crystal structures and predicts material properties from high‑resolution microscopy images. Deployed as an end‑to‑end pipeline integrating denoising, physics‑guided template retrieval, symmetry‑constrained reconstruction, and ML‑based property prediction, it bridges experimental STEM imaging with atomistic simulation—achieving accurate, closed‑loop reasoning from microscopy to materials modeling.